OpenStudy (anonymous):

how do i find the lcd for large denominators like19/70-19/110.

OpenStudy (anonymous):

look at the prime factors of 70 and 110: 70 = 2*5*7 110 = 2*5*11 to get the LCD, take the factors they don't have in common 7 and 11 and multiply that with what they have in common, 2*5. So your answer should be (2*5)*7*11=770.

OpenStudy (anonymous):

sorry, not your answer but the lcd to the problem...

OpenStudy (mertsj):

Write the prime factors in exponential form. Select the largest power of every DIFFERENT factor. That will be your LCM

OpenStudy (mertsj):